# On extension modules of finite homological dimension

### v1

@@ 4. Finite Homological dimensions of deficiency modules
From now until the end of the text, $R$ is local, and unless otherwise stated, all $R$-modules will be finitely generated.
In this section, we derive some results about the implication of having modules of deficiency with certain finite homological dimensions. For this, we assume here that $R$ has a dualizing complex $\boldsymbol{D}$. Moreover, we suppose that $\boldsymbol{D}$ is normalized in the sense that $\operatorname{Hom}_{R}(k,\boldsymbol{D})\simeq\Sigma^{d}k$, where $d$ is the dimension of $R$.
The existence of $\boldsymbol{D}$ is equivalent to $R$ being a factor of a Gorenstein local ring $S$, i.e., $R$ is an epimorphic image of $S$. In that context, Schenzel [33] generalized the notion of a canonical module of a ring in the following sense: For an $R$-module $M$, set
@@
for all $i\in\mathbb{Z}$. For each $i$, the $R$-module $K^{i}(M)$ is called the $i$-th deficiency module of $M$. The $R$-module $K(M):=K^{\dim_{R}(M)}(M)$ is called the canonical module of $M$. The modules of deficiency satisfy the following isomorphisms:
$H^{i}_{\mathfrak{m}}(M)\cong\text{Hom}_{R}(K^{i}(M),E_{R})\text{ and }K^{i}(M)\cong\operatorname{H}_{i-d}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D})),$
where $E_{R}$ is the hull envelope of the residue field of $R$. Note that the first isomorphism shows that $K^{i}(M)=0$ for $i<0pt(M)$ or $i>\dim(M)$, and $K^{i}(M)\neq 0$ for $i=0pt(M)$ and $i=\dim(M)$. For further details about deficiency modules, we refer the reader to [33, 34].
In [13, 17], the authors investigated the consequences of the finiteness of classical homological dimensions, namely, projective and injective dimensions, of the deficiency modules and the canonical module. Motivated by their work, we apply Theorem 3.6 to extend those results to a broader class of homological dimensions considered in this paper. We also emphasize that our approach differs from theirs in both method and perspective.
@@
For each $i$, we note that $K^{i}(M)=\operatorname{H}_{i-d}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D}))=\operatorname{Ext}_{R}^{d-i}(M,\boldsymbol{D})$. Since $K^{i}(M)=0$ for $i<0pt(M)$ or $i>\operatorname{dim}_{R}(M)$, by assumption we see have that all modules $\operatorname{Ext}_{R}^{i}(M,\boldsymbol{D})$ have finite $\operatorname{H-dim}_{R}$ or all modules have finite $\operatorname{Hid}$. Then the result follows from [37, Proposition 4.2] and Theorem 3.6. ∎
The following corollary contains the counterparts of [14, Proposition 3.5(b)] with Gorenstein dimensions.
@@
In view of [33, Theorem 1.14], it is enough to prove (1). One implication follows from Theorem 4.1(1). For the other, suppose that $\operatorname{Hid}_{R}(M)<\infty$. Since $M\cong K(K(M))$, then $K(M)$ is Cohen-Macaulay ([33, Theorem 1.14] again) and from Theorem 4.1(2) it follows that $\operatorname{H-dim}_{R}(K(M))<\infty$. ∎
5. On Finite Homological dimension of Ext modules over CM rings
@@
Proposition 5.1 (Proposition 2.4 & [6, Theorem 3.4.11]).
@@
    $\operatorname{G-dim}_{R}(M)<\infty$ if and only if $M\in\mathcal{A}_{\omega}(R)$.
@@
    $\operatorname{Gid}_{R}(M)<\infty$ if and only if $M\in\mathcal{B}_{\omega}(R)$.
@@
In [12, Corollary 42(3)], although the authors provide an independent proof of Proposition 5.1, this proposition assures that their vanishing hypotheses are not necessary.
@@
Suppose that $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$. In particular, $\operatorname{G-dim}(\operatorname{Hom}_{R}(M,N))<\infty$. Note that $\operatorname{Hom}_{R}(M,N)$ and $N$ belong to $\mathcal{A}_{\omega}(R)$. Then by [27, Lemma 5.1], the map $\Phi_{M,N}$ is an isomorphism. Hence, it follows from Proposition 5.1(3) that $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(M,N\otimes_{R}\omega))<\infty$.
@@
where the last isomorphism is because $N\in\mathcal{A}_{\omega}(R).$ Thus, $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$. ∎
Let $\mathcal{A}$ be an abelian category and let $\mathcal{X}$ a class of objects in $\mathcal{A}$. We say that $\mathcal{X}$ is closed under short exact sequences if, in each short exact sequence, the fact that two objects are in $\mathcal{X}$ implies that the third so is. We say that $\mathcal{X}$ is closed under isomorphisms if every object isomorphic to $\mathcal{X}$, is in $\mathcal{X}$.
@@
We must show that for all $0\leq j\leq n$, we have $\operatorname{Hom}_{R}(\Omega^{j}(M),N))\in\mathcal{X}$. We prove this by induction on $j$. For $j=0$, this holds by hypothesis. Now suppose $0<j\leq n$ and that $\operatorname{Hom}_{R}(\Omega^{j-1}(M),N))<\infty$. We have an exact sequence
[…]
which yields
[…]
since $\operatorname{Ext}^{1}_{R}(\Omega^{j-1}(M),N)\simeq\operatorname{Ext}^{j}_{R}(M,N)$. Such an exact sequence splits into
[…]
and
[…]
Since $\mathcal{X}$ is closed under short exact sequences and isomorphisms, it follows that $\operatorname{Hom}_{R}(\Omega^{j}(M),N))\in\mathcal{X}$. ∎
The following theorem improves [27, Lemma 5.1].
@@
For $i=0$, the isomorphism holds from [27, Lemma 5.1]. For $i>0$, consider the short exact sequence
[…]
@@
[…]
where the second row is exact. As to the first one, since $\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N)\cong\operatorname{Ext}^{i}_{R}(M,N)$, by assumption, [35, 1.9(a)] and Lemma 5.4, all $R$-modules of this exact sequence belong to $\mathcal{A}_{C}(R)$. Therefore, the first row turns out to be exact as well by [31, Lemma 3.1.12].
Additionally, by [27, Lemma 5.1], the vertical maps are isomorphisms. Then, from the five lemma, we assure the existence of a natural isomorphism $\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N)\otimes_{R}C\cong\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}C)$, from which we conclude the result. ∎
@@
Assume $R$ Cohen-Macaulay with canonical module $\omega$, and let $n$ be a non-negative integer. Let $M$ and $N$ be $R$-modules such that $\operatorname{H-dim}_{R}(N)<\infty$. Then the following conditions are equivalents:
@@
We first prove the equivalence. Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ (resp. $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$) for all $0\leq i\leq n$. Given $0\leq i\leq n$, we must show that $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$ (resp. $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$). The case $i=0$ follows from Proposition 5.3. Now, suppose $i>0$. Consider the exact sequence
$0\stackrel{{\scriptstyle~{}}}{{\longrightarrow}}\Omega^{i}(M)\stackrel{{\scriptstyle~{}}}{{\longrightarrow}}R^{r_{i}}\stackrel{{\scriptstyle~{}}}{{\longrightarrow}}\Omega^{i-1}(M)\longrightarrow 0$
Then it induces an exact sequence
(5.1) $\scriptstyle{0\longrightarrow\operatorname{Hom}_{R}(\Omega^{i-1}(M),N\otimes_{R}\omega)\longrightarrow\operatorname{Hom}_{R}(R^{r_{i}},N\otimes_{R}\omega)\longrightarrow\operatorname{Hom}_{R}(\Omega^{i}(M),N\otimes_{R}\omega)\longrightarrow\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega)\longrightarrow 0}.$
(resp.
(5.2) $\scriptstyle{0\longrightarrow\operatorname{Hom}_{R}(\Omega^{i-1}(M),N)\longrightarrow\operatorname{Hom}_{R}(R^{r_{i}},N)\longrightarrow\operatorname{Hom}_{R}(\Omega^{i}(M),N)\longrightarrow\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N)\longrightarrow 0}\mbox{)}.$
By Proposition 5.1(4), $\operatorname{Hid}_{R}(N\otimes_{R}\omega)<\infty$. Additionally, by Lemma 5.4 and Theorem 2.1, $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N))<\infty$ (resp. $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N\otimes_{R}\omega))<\infty$) for $j=i-1,i$, so by Proposition 5.3, $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N\otimes_{R}\omega))<\infty$ (resp. $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N))<\infty$) for $j=i-1,i$. Thus, the first three terms of the exact sequence (5.1) (resp. (5.2)) have finite $\operatorname{Hid},$ (resp. $\operatorname{H-dim}_{R}$) and therefore $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega))<\infty$ (resp. $\operatorname{H-dim}_{R}\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N)<\infty$). The result follows from the fact that $\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega)\cong\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega)$.
Now, we prove the last statement. Suppose that (1) holds. Then the $R$-modules $N$ and $\operatorname{Ext}_{R}^{i}(M,N)$, for all $i=0,\ldots,n$, have finite Gorenstein dimension. Thus, these modules belong to $\mathcal{A}_{\omega}(R)$. Hence, the isomorphism required follows from Theorem 5.5. ∎
@@
Assume $R$ is Cohen-Macaulay, and let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(N)<\infty$. If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then $\operatorname{H-dim}_{R}(M)<\infty$.
@@
We may assume that $R$ is complete, and hence $R$ has a canonical module $\omega$. Since $\operatorname{pd}_{R}(N)<\infty$, by Proposition 5.1, we have $\operatorname{id}_{R}(N\otimes_{R}\omega)<\infty$ and, by Theorem 5.6, $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$ for all $0\leq i\leq t-0pt(M)$. Then it follows from Corollary 3.7 that $\operatorname{H-dim}_{R}(M)<\infty$. ∎
The following corollary contains a generalization of [12, Corollary 30] and [28, Theorem 4.2$(3)\Rightarrow(2)$].
Corollary 5.8.
Assume $R$ is Cohen-Macaulay, and let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(N)<\infty$ and $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$. If $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $1\leq i\leq t-0pt(M)$, then $\operatorname{H-dim}_{R}(M)=0$.
Proof.
By Corollary 5.7, we have that $\operatorname{H-dim}_{R}(M)<\infty$. As $\operatorname{pd}_{R}(N)<\infty$, we have that $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $i>t-0ptM$ by [1, Theorem 4.13], and that $\operatorname{H-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,N)\not=0\}$ by [30, Theorem 2.6(ii)]. Therefore, $\operatorname{H-dim}_{R}(M)=0$. ∎
Combining Proposition 3.4 with Corollary 5.7, we have the following byproduct.
@@
Next, we explore the consequences of Theorem 5.6 when in such theorem we consider $N=R$. Recall that the type of an $R$-module $M$ is defined to be the dimension of the $k$-vector space $\operatorname{Ext}^{r}_{R}(k,M)$, where $r=0pt(M)$, and denoted by $\operatorname{r}(M)$.
@@
Now, by taking the tensor product by the residue field of $R$ and dimensions, the required equality holds from [5, Proposition 3.3.11(c)]. ∎
@@
Assume $R$ is Cohen-Macaulay, and let $M$ be an $R$-module of depth $r$ and dimension $s$. Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $t-s\leq i\leq t-r$ Then:
  1. (1)
    $\operatorname{H-dim}_{R}(M)<\infty$.
  2. (2)
    $\operatorname{dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))\leq t-i$ for all $i\neq t-s$ and $\dim_{R}(\operatorname{Ext}^{t-s}_{R}(M,R))=s$.
  3. (3)
    $\operatorname{r}(M)=\mu(\operatorname{Ext}^{t-r}_{R}(M,R))\operatorname{r}(R).$
  4. (4)
    If $M$ has type one, then $R$ is Gorenstein.

### v2

@@ 3. Finite Homological dimensions of deficiency modules
In the rest of the paper, $R$ is local, and unless otherwise stated, all $R$-modules will be finitely generated. Moreover, in this section, we assume that $R$ has a dualizing complex $\boldsymbol{D}$. Also, we assume that $\boldsymbol{D}$ is normalized in the sense that $\sup D=\dim R$.
The existence of $\boldsymbol{D}$ is equivalent to $R$ being a factor of a Gorenstein local ring $S$, i.e., $R$ is a homomorphic image of $S$. In that context, Schenzel [29] generalized the notion of a canonical module of a ring in the following sense: For an $R$-module $M$, set
@@
for all $i\in\mathbb{Z}$. For each $i$, the $R$-module $K^{i}(M)$ is called the $i$-th deficiency module of $M$. The $R$-module $K(M):=K^{\dim_{R}(M)}(M)$ is called the canonical module of $M$. By local duality we have $K^{i}(M)=0$ for $i<0pt(M)$ or $i>\dim(M)$, and $K^{i}(M)\neq 0$ for $i=0pt(M)$ and $i=\dim(M)$. The modules of deficiency also satisfy the following isomorphism for all $i$:
$K^{i}(M)\cong\operatorname{H}_{i}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D})).$
For further details about deficiency modules, we refer the reader to [29].
In [12], the authors investigated the consequences of the finiteness of classical homological dimensions, namely, projective and injective dimensions, of the deficiency modules. Motivated by their work, we apply Theorem 2.5 to extend those results to Gorenstein dimensions. We also emphasize that our approach differs from theirs in both method and perspective.
@@
For each $i$, we note that $K^{i}(M)=\operatorname{H}_{i}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D}))=\operatorname{Ext}_{R}^{-i}(M,\boldsymbol{D})$. Since $K^{i}(M)=0$ for $i<0pt(M)$ or $i>\operatorname{dim}_{R}(M)$, by assumption we see have that all modules $\operatorname{Ext}_{R}^{i}(M,\boldsymbol{D})$ have finite $\operatorname{H-dim}_{R}$ or all modules have finite $\operatorname{Hid}$. Then the result follows from [32, Proposition 4.2] and Theorem 2.5. ∎
The following corollary contains the counterparts of [13, Proposition 3.5(b)] to Gorenstein dimensions.
@@
In view of [29, Theorem 1.14], it is enough to prove (1). One implication follows from Theorem 3.1(1). For the other, suppose that $\operatorname{Hid}_{R}(M)<\infty$. Since $M\cong K(K(M))$, then $K(M)$ is Cohen-Macaulay ([29, Theorem 1.14] again) and from Theorem 3.1(2) it follows that $\operatorname{H-dim}_{R}(K(M))<\infty$. ∎
Corollary 3.3.
Let $R$ be a Cohen-Macaulay local ring with canonical module $\omega$. Assume that $\operatorname{Ext}_{R}^{i}(\omega,R)=0$ for all $1\leq i\leq t$. Let $^{\dagger}$ stand for $\operatorname{Hom}_{R}(\_,\omega)$. If $\operatorname{Gid}_{R}((\omega^{*})^{\dagger})<\infty$, then $R$ is Gorenstein.
Proof.
Note that from the vanishing hypothesis and [18, Theorem 2.1(i)] we see that $\omega^{*}$ is maximal Cohen-Macaulay. Since $\operatorname{Gid}_{R}((\omega^{*})^{\dagger})<\infty$, then by Corollary 3.2, $\operatorname{G-dim}_{R}(\omega^{\ast})<\infty$. Therefore, the conclusion follows from [24, Corollary 5.13]. ∎
4. A duality for Ext modules with finite homological dimension
@@
Proposition 4.1 ([9, Theorem 4.4], [6, Corollary 3.1.11 & Theorem 3.4.11] ).
@@
    $\operatorname{G-dim}_{R}(M)<\infty$ if and only if $M\in\mathcal{A}(R)$.
@@
    $\operatorname{Gid}_{R}(M)<\infty$ if and only if $M\in\mathcal{B}(R)$.
@@
In [11, Corollary 42(3)], although the authors provide this without using Proposition 4.1, this proposition assures that their vanishing assumptions are not necessary.
@@
Suppose that $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$. In particular, $\operatorname{G-dim}(\operatorname{Hom}_{R}(M,N))<\infty$. Note that $\operatorname{Hom}_{R}(M,N)$ and $N$ belong to $\mathcal{A}(R)$. Then by [23, Lemma 5.1], the map $\Phi_{M,N}$ is an isomorphism. Hence, it follows from Proposition 4.1(3) that $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(M,N\otimes_{R}\omega))<\infty$.
@@
where the last isomorphism follows since $N\in\mathcal{A}(R).$ Thus, $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$. ∎
Let $\mathcal{A}$ be an abelian category and let $\mathcal{X}$ be a class of objects in $\mathcal{A}$. We say that $\mathcal{X}$ is closed under short exact sequences if the two out of three property holds for short exact sequences of objects in $\mathcal{X}$. We say that $\mathcal{X}$ is closed under isomorphisms if every object isomorphic to an object in $\mathcal{X}$ is in $\mathcal{X}$.
@@
We must show that for all $0\leq j\leq n$, we have $\operatorname{Hom}_{R}(\Omega^{j}(M),N))\in\mathcal{X}$. We prove this by induction on $j$. For $j=0$, this holds by hypothesis. Now suppose $0<j\leq n$ and that $\operatorname{Hom}_{R}(\Omega^{j-1}(M),N))\in\mathcal{X}$. We have an exact sequence ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Omega^{j}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{n_{j}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Omega^{j-1}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$ which yields
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}_{R}(\Omega^{j-1}(M),N)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{n_{j}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}_{R}(\Omega^{j}(M),N)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Ext}^{j}_{R}(M,N)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
since $\operatorname{Ext}^{1}_{R}(\Omega^{j-1}(M),N)\simeq\operatorname{Ext}^{j}_{R}(M,N)$. The result follows by breaking this exact sequence into two short ones and using the fact that $\mathcal{X}$ is closed under short exact sequences and isomorphisms.
The following theorem improves [23, Lemma 5.1]. For nomenclature on semidualizing modules and their corresponding Auslander classes, we refer to [27].
@@
For $i=0$, the isomorphism holds from [23, Lemma 5.1]. For $i>0$, consider the short exact sequence ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Omega^{i}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{r_{i}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Omega^{i-1}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
@@
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}_{R}(\Omega^{i-1}(M),N)\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\Omega^{i-1}(M),N}}$${\operatorname{Hom}_{R}(R^{r_{i}},N)\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{R^{r_{i}},N}}$${\operatorname{Hom}_{R}(\Omega^{i}(M),N)\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\Omega^{i}(M),N}}$${\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N)\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}(\Omega^{i-1}(M),N\otimes_{R}C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}(R^{r_{i}},N\otimes_{R}C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}(\Omega^{i}(M),N\otimes_{R}C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N\otimes_{R}C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
where the second row is exact. As to the first one, since $\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N)\cong\operatorname{Ext}^{i}_{R}(M,N)$, by assumption, [30, 1.9(a)] and Lemma 4.4, all $R$-modules of this exact sequence belong to $\mathcal{A}_{C}(R)$. Therefore, the first row turns out to be exact as well by [27, Lemma 3.1.12].
Additionally, by [23, Lemma 5.1], the vertical maps are isomorphisms. The five lemma settles the desired isomorphism. ∎
Recall that the classes of modules of finite projective dimension, finite injective dimension, finite Gorenstein dimension, and finite Gorenstein injective dimension are closed under short exact sequences; see [10, Corollaries 8.1.9 and 8.2.9, Proposition 9.2.15] and [6, Corollary 1.2.9]. With these facts in mind, together with the results established earlier in this section, we are now able to prove our main theorem.
@@
Assume $R$ is Cohen-Macaulay with canonical module $\omega$, and let $n$ be a non-negative integer. Let $M$ and $N$ be $R$-modules such that $\operatorname{H-dim}_{R}(N)<\infty$. Then the following conditions are equivalent:
@@
Once the equivalence is proved, all the isomorphisms follow from Theorem 4.5, Proposition 4.1, and item (1).
Now, we prove $(1)\Rightarrow(2)$. Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq n$. Given $0\leq i\leq n$, we must show that $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$. The case $i=0$ follows from Proposition 4.3. Now, suppose $i>0$. Consider the exact sequence $0\stackrel{{\scriptstyle~}}{{\longrightarrow}}\Omega^{i}(M)\stackrel{{\scriptstyle~}}{{\longrightarrow}}R^{r_{i}}\stackrel{{\scriptstyle~}}{{\longrightarrow}}\Omega^{i-1}(M)\longrightarrow 0.$ Then it induces an exact sequence
$\scriptscriptstyle{0\longrightarrow\operatorname{Hom}_{R}(\Omega^{i-1}(M),N\otimes_{R}\omega)\longrightarrow\operatorname{Hom}_{R}(R^{r_{i}},N\otimes_{R}\omega)\longrightarrow\operatorname{Hom}_{R}(\Omega^{i}(M),N\otimes_{R}\omega)\longrightarrow\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega)\longrightarrow 0}.$
By Proposition 4.1(4), $\operatorname{Hid}_{R}(N\otimes_{R}\omega)<\infty$. Additionally, by Lemma 4.4, $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N))<\infty$ for j=i1,ij=i-1,i, so by Proposition 4.3, $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N\otimes_{R}\omega))<\infty$ for j=i1,ij=i-1,i. Thus, the first three terms of the exact sequence above have finite $\operatorname{Hid},$ and therefore $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega))<\infty$. The result follows from the fact that $\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega)\cong\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega)$.
As to $(2)\Rightarrow(1)$, we follow the same lines as before by considering the exact sequence
$\scriptstyle{0\longrightarrow\operatorname{Hom}_{R}(\Omega^{i-1}(M),N)\longrightarrow\operatorname{Hom}_{R}(R^{r_{i}},N)\longrightarrow\operatorname{Hom}_{R}(\Omega^{i}(M),N)\longrightarrow\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N)\longrightarrow 0}$
for $0<i\leq n$. ∎
[…]
@@
Assume $R$ is Cohen-Macaulay local, and let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(N)<\infty$. If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then $\operatorname{H-dim}_{R}(M)<\infty$. In addition, if $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $1\leq i\leq t-0pt(M)$, then $\operatorname{H-dim}_{R}(M)=0$.
@@
We may assume that $R$ is complete, and hence $R$ has a canonical module $\omega$. Since $\operatorname{pd}_{R}(N)<\infty$, by Proposition 4.1, we have $\operatorname{id}_{R}(N\otimes_{R}\omega)<\infty$ and, by Theorem 4.6, $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$ for all $0\leq i\leq t-0pt(M)$. Then it follows from Corollary 2.6 that $\operatorname{H-dim}_{R}(M)<\infty$.
Now, by the vanishing assumption, as $\operatorname{pd}_{R}(N)<\infty$, we have that $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all i>t0pt(M)i>t-0pt(M) by [1, Theorem 4.13], and that $\operatorname{H-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,N)\not=0\}$ by [26, Theorem 2.6(ii)]. Therefore, $\operatorname{H-dim}_{R}(M)=0$. ∎
Combining Corollary 2.3 with Corollary 4.7, we have the following byproduct.
@@
Next, we explore the consequences of Theorem 4.6 when one considers $N=R$. Recall that the type of an $R$-module $M$ is defined to be the dimension of the $k$-vector space $\operatorname{Ext}^{r}_{R}(k,M)$, where $r=0pt(M)$, and denoted by $\operatorname{r}(M)$.
@@
Now, by taking the tensor product by the residue field of $R$ and counting dimensions, the required equality holds from [5, Proposition 3.3.11(c)]. ∎
@@
Assume $R$ is Cohen-Macaulay, and let $M$ be an $R$-module of depth $r$ and dimension $s$. Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $t-s\leq i\leq t-r$. Then $\operatorname{H-dim}_{R}(M)<\infty$. Moreover, $\operatorname{dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))\leq t-i$ for all $i\neq t-s$ and $\dim_{R}(\operatorname{Ext}^{t-s}_{R}(M,R))=s$.